PHOTOELECTRON-PHOTOFRAGMENT COINCIDENCE SPECTROSCOPY OF TERT-BUTOXIDE AND THE CARBANION ISOMER

\underline{BEN B. SHEN}, BERWYCK L. J. POAD, AMELIA W. RAY, and ROBERT E. CONTINETTI^{a}, Department of Chemistry and Biochemistry, University of California, San Diego, CA 92093.

The photodetachment dynamics of the C_{4}H_{9}O^{−} anion, produced by proton abstraction from tert-butanol has been studied using a cryogenically-cooled linear electrostatic ion trap to store a fast ion beam^{b}. The ion trap is coupled with a Photoelectron-Photofragment Coincidence spectrometer for a kinematically complete experiment in which dissociative neutral intermediates of known internal energy are created by photodetachment of a stable molecular anion. By collecting all resultant neutral fragments from the dissociation event in coincidence with the detached electron using time- and position-sensitive detectors, the correlated momenta and energies of all products can be determined. At 388, 537, and 600nm, both stable and dissociative photodeatchment processes of C_{4}H_{9}O^{−} are observed. The stable channel is consistent with previous photodetachment spectrum from tert-butoxide anion^{c} as well as a Franck-Condon simulation of the photoelectron spectrum. The dissociative channel is consistent with a methyl loss from an alkylhydroxy radical. Results will be analyzed in the context of calculations employing MP2/6-311++G(d,p) and CBSQ level of theory.

aThis work is supported by the United States Department of Energy under grant number DE-FG03-98ER14879

bC. J. Johnson, B. B. Shen, B. L. J. Poad, and R. E. Continetti, Rev. Sci. Instrum., 82, 105105 (2011)

cT. M. Ramond, G. E. Davico, R. L. Schwartz, W. C. Lineberger, J. Chem. Phys., 112, 1158 (2000)